# PIS3R: Very Large Parallax Image Stitching via Deep 3D Reconstruction

### v1

@@ 3.5 Implementation
$\bullet$ Diffusion model. To train the RDDM model, we created paired training data consisting of degraded and corresponding ground truth images. To better suit our task, we manually processed ground truth images from the RainDrop dataset  (Qian et al. 2018) to generate the customized 3DIS-D dataset, incorporating various noises for enhanced robustness beyond reprojection artifacts and occlusions.
We adopted a progressive training strategy with a global learning rate of $8e^{-5}$ and batch size of $1$. In the first stage, we trained on 3DIS-D at $512\times 512$ resolution with $10$ sampling steps for $120,000$ iterations, using ground truth as targets and processed images as degraded inputs. In the second stage, we fine-tuned on GoPro (Nah, Hyun Kim, and Mu Lee 2017) dataset at the same resolution with $5$ sampling steps for $20,000$ iterations. Notably, our approach achieves satisfactory results with limited training data. All training and testing were performed on an NVIDIA RTX $4090$ GPU.
4 Discussion
$\bullet$ Is PIS3R suitable for pure rotation or slight parallax? PIS3R delivers comparable visual quality to state-of-the-art methods under pure rotation or slight parallax (Fig. 3, Appendix) but requires more processing time per image than UDIS++ (Tab. 1, Appendix). We recommend PIS3R for scenarios with large parallax and propose a parallax classifier (Sec. B, Appendix) to estimate input image parallax levels.
$\bullet$ Why deep 3DR? Deep 3DR solutions such as VGGT demonstrate significantly greater robustness and accuracy than classical SfM like COLMAP in estimating camera intrinsic and extrinsic parameters, particularly in challenging two-view scenarios such as pure rotation and very large parallax (as shown in Fig. 1 and detailed in the Sup material). Moreover, the point maps predicted by VGGT offer substantially higher density compared to traditional SfM point clouds, providing strong support for the following image stitching reprojection step.
$\bullet$ Why PIS3R is geometry-preserve? PIS3R employs a deep-based reconstruction framework to accurately predict the point cloud representation and the camera intrinsic and extrinsic parameters. Accurate pose estimation of the images is fundamental to preserving geometric structure in PIS3R. This is also the reason why SOTA methods fail to maintain geometric structure.As illustrated in Sec. 5.3 and tab. 2.
$\bullet$ How does PIS3R support 3D vision downstream tasks? Our preliminary stitched image is generated via reprojection based on the camera parameters. This provides the stitched image with accurate geometric spatial information. As shown in Fig. 6 and Sec. 5.3, stitching results of PIS3R in very large parallax scenarios can be directly input into a geometry-based 3D reconstruction framework for point cloud reconstruction, while better maintaining the geometric structure and integrity of the scene due to its excellent geometry preservation capability.
@@ 4 Discussion
Figure 6: The results of MetaShape reconstruction using stitched images. When used as inputs for 3D reconstruction in MetaShape, images stitched with APAP and ELA resulted in large number of surface holes. Images stitched with UDIS++ exhibited significant distortions to geometric structure. In contrast, reconstructions generated from our stitched images consistently exhibited high geometric consistency and completeness within MetaShape. This enhanced structural integrity suggests that PIS3R is more suitable for downstream 3D reconstruction tasks.
5 Experiments
Figure 7: Qualitative comparison of stitched images on synthetic dataset and real-world dataset. Our PIS3R generates images with significantly more complete structures and fewer severe distortions compared to existing image stitching algorithms APAP, ELA and UDIS++ in both datasets. Ref./Tar. Img. refers to Reference and Target Images.
@@ 5 Experiments
In this section, we conduct a comprehensive series of comparative experiments to evaluate the performance of our method PIS3R. We compare PIS3R against three state-of-the-art image stitching solutions: APAP (Zaragoza et al. 2014), ELA (Li et al. 2018) and UDIS++ (Nie et al. 2023).
5.1 Datasets
@@
$\bullet$ Synthetic Dataset: As shown in Fig. 8a, this test dataset was rendered using Blender. It comprises 20 distinct scenes  [^3] https://sketchfab.com  [^4] https://blenderco.cn/ 555 https://free3d.com/zh/3d-models/blender with 600 image pairs and their corresponding ground truth (GT) images. Most image pairs have large-baseline, large-parallax characteristics.
@@
$\bullet$ Real-World Dataset: As shown in Fig. 8b, we capture the dataset with a mobile phone. It comprises 10 distinct scenes with 682 image pairs and their corresponding GT images. We test this real-world dataset to validate the performance of PIS3R in real-world scenarios.
@@
Figure 8: Sample images from the experimental dataset. (a)Synthetic dataset rendered by Blender (Community 2018). (b)Real-world dataset captured by ourselves. Additionally, each scene image and its ground-truth image share the same camera extrinsic.
5.2 Comparison of Image Stitching
Evaluation Metrics.
@@ 5.2 Comparison of Image Stitching
The traditional image stitching paradigm consists of warping and composition stages. However, as a 3D-reconstruction-based image stitching paradigm, PIS3R does not include these steps. Therefore, overlapping regions based metrics of warping results is not suitable for PIS3R. Notably, each reference image shares the same camera extrinsics with its GT image (Fig. 8), and scene images are obtained by center-cropping GT images. Theoretically, stitched images should share the same extrinsics with reference images. Thus, we propose a novel three-step metric computation approach (Fig.4 in Appendix):
@@
We employ metrics including: PSNR, SSIM  (Wang et al. 2004), LPIPS  (Zhang et al. 2018). Furthermore, to avoid metric inaccuracies caused by registration failures, we calculated the Registration Success Rate (RSR):
@@
Image Stitching in Synthetic Dataset.
@@
$\bullet$ Qualitative Comparison. The qualitative comparison results are shown in Fig. 7a, where the reference and target images are displayed in the left-most column. APAP, ELA and UDIS++ exhibit noticeable artifacts and geometric inconsistencies, while Our results better preserve the geometric structure of the scene. Additional cases are provided in Fig. 1 of Appendix.
$\bullet$ Quantitative Comparison. The quantitative comparison results are presented in Tab. 1. Our PIS3R outperforms the other three methods across all metrics when including all data in the calculations. We still maintain a substantial lead in PSNR, SSIM after excluding failed registration cases. What is noteworthy is that RSR substantially indicates the method’s ability to maintain geometric consistency.
All Image Pairs W/O Failed Registration Cases
APAP 19.116 0.659 0.169 87.5% 20.063 0.691 0.146
Synthetic ELA 19.813 0.67 0.158 91.5% 20.487 0.719 0.139
Dataset UDIS++ 19.858 0.690 0.139 86.2% 21.038 0.748 0.097
PIS3R 22.026 0.782 0.110 95.7% 22.351 0.796 0.102
APAP 16.764 0.525 0.246 81.6% 18.047 0.572 0.207
Real World ELA 16.983 0.547 0.245 98.2% 17.064 0.553 0.242
Dataset UDIS++ 17.390 0.554 0.204 90.3% 18.874 0.615 0.208
PIS3R 19.674 0.644 0.180 100.0% 19.674 0.644 0.180
Table 1: Quantitative comparison of image stitching on synthetic dataset and real world dataset. PIS3R outperforms the other three methods across PSNR, SSIM and LPIPS when all image pairs are included in the evaluation. Moreover, even after excluding registration-failed cases, PIS3R still maintains a noticeable advantage in PSNR and SSIM.
Image Stitching in Real-World Dataset.
@@
$\bullet$ Qualitative Comparison. The qualitative comparison results are presented in Fig. 7b. APAP, ELA and UDIS++ exhibit noticeable artifacts and geometric inconsistencies. In contrast, our PIS3R method consistently generates images with significantly more complete structures and fewer severe distortions in these complex and challenging scenarios. Additional cases are provided in Fig. 2 of Appendix.
$\bullet$ Quantitative Comparison. As shown in Tab. 1, PIS3R demonstrates superior performance across all evaluation metrics on the complete real-world dataset. Moreover, after excluding failed registration cases, it continues to achieve leading performance on all metrics.
5.3 Geometric Consistency Verification of Stitched Images
@@
[…]
$\bullet$ Quantitative Comparison. As shown in Table 2, PIS3R achieves a significantly lower epipolar error for stitched images exhibiting very large parallax conditions, notably outperforming the other three approaches.
APAP ELA UDIS++ PIS3R
$d_{\text{Sam}}^{2}$(pixel) $\downarrow$ 22.78 38.97 12.94 4.12
Table 2: Sampson epipolar error of stitching image on synthetic dataset. PIS3R achieves a lower epipolar error for stitched images exhibiting very large parallax conditions.
5.4 Ablation studies
Comparison of Different Diffusion Models.
As shown in Fig. 5, we compared three image restoration diffusion models: DiffIR  (Xia et al. 2023) [^6] https://github.com/Zj-BinXia/DiffIR, DDRM (Kawar et al. 2022) [^7] https://github.com/bahjat-kawar/ddrm, and RDDM  (Liu et al. 2024) [^8] https://github.com/nachifur/RDDM. We train DiffIR and RDDM on 3DIS-D dataset(as illustrated in Sec. 3.5) and present comparative results of the restoration performance. Notably, DDRM does not rely on task-specific training and can be applied directly. RDDM can effectively remove noise while preserving geometric structures.
Quantitative Impact of Diffusion Models.
We quantify the metric differences between reprojected images processed and unprocessed by the diffusion model. As shown in Tab.2,3 in Appendix, diffusion model notably enhances stitched images’ visual quality while having a negligible impact on the evaluation metrics. This trade-off is thus deemed both acceptable and necessary.
6 Conclusion

### v2

@@ III-E Implementation
$\bullet$ Diffusion model. To train the RDDM model, we created paired training data consisting of degraded and corresponding ground truth images. To better suit our task, we manually processed ground truth images from the RainDrop dataset  [raindrop_qian2018attentive] to generate the customized 3DIS-D dataset, incorporating various noises for enhanced robustness beyond reprojection artifacts and occlusions.
We adopted a progressive training strategy with a global learning rate of $8e^{-5}$ and batch size of $1$. In the first stage, we trained on 3DIS-D at $512\times 512$ resolution with $10$ sampling steps for $120,000$ iterations, using ground truth as targets and processed images as degraded inputs. In the second stage, we fine-tuned on GoPro [gopro_nah2017deep] dataset at the same resolution with $5$ sampling steps for $20,000$ iterations. Notably, our approach achieves satisfactory results with limited training data. All training and testing were performed on an NVIDIA RTX $4090$ GPU.
IV Discussion
$\bullet$ Is PIS3R suitable for pure rotation or slight parallax? PIS3R delivers comparable visual quality to state-of-the-art methods under pure rotation or slight parallax (Fig. 10) but requires more processing time per image than UDIS++ (Tab. VII). We recommend PIS3R for scenarios with large parallax and propose a parallax classifier (Sec. B, Appendix) to estimate input image parallax levels.
$\bullet$ Why deep 3DR? Deep 3DR solutions such as VGGT demonstrate significantly greater robustness and accuracy than classical SfM like COLMAP in estimating camera intrinsic and extrinsic parameters, particularly in challenging two-view scenarios such as pure rotation and VLP (as shown in Fig. 1 and detailed in the Sup material). Moreover, the point maps predicted by VGGT offer substantially higher density compared to traditional SfM point clouds, providing strong support for the following image stitching reprojection step.
$\bullet$ Why PIS3R is geometry-preserve? PIS3R employs a deep-based reconstruction framework to accurately predict the point cloud representation and the camera intrinsic and extrinsic parameters. Accurate pose estimation of the images is fundamental to preserving geometric structure in PIS3R. This is also the reason why SOTA methods fail to maintain geometric structure.As illustrated in Sec. V-C and tab. IV.
$\bullet$ How does PIS3R support 3D vision downstream tasks? Our preliminary stitched image is generated via reprojection based on the camera parameters. This provides the stitched image with accurate geometric spatial information. As shown in Fig. 6 and Sec. V-C, stitching results of PIS3R in VLP scenarios can be directly input into a geometry-based 3D reconstruction framework for point cloud reconstruction, while better maintaining the geometric structure and integrity of the scene due to its excellent geometry preservation capability.
@@ IV Discussion
Figure 6: The results of MetaShape reconstruction using stitched images. When used as inputs for 3D reconstruction in MetaShape, images stitched with APAP and ELA resulted in large number of surface holes. Images stitched with UDIS++ and OBJ-GSP exhibited significant distortions to geometric structure. In contrast, reconstructions generated from our stitched images consistently exhibited high geometric consistency and completeness within MetaShape. This enhanced structural integrity suggests that PIS3R is more suitable for downstream 3D reconstruction tasks.
V Experiments
@@
In this section, we conduct a comprehensive series of comparative experiments to evaluate the performance of our method PIS3R. We compare PIS3R against three state-of-the-art image stitching solutions: APAP [apap_zaragoza2014projective], ELA [ELA_li2018parallax], UDIS++ [udis2_nie2023parallax] and OBJ-GSP [obj-gsp_cai2025object].
V-A Datasets
PIS3R proposes a solution for image stitching under very large parallax(VLP)—a scenario not adequately addressed by existing datasets. To evaluate its effectiveness, we introduce a self-collected large-parallax dataset comprising both synthetic and real image pairs. Furthermore, to demonstrate its applicability under slight parallax, we also evaluate our method on the UDIS-D  [udis_d_nie2021unsupervised] dataset.
V-A1 UDIS-D Dataset
UDIS-D is a large-scale image dataset designed for image stitching or image registration. It includes image pairs with varying overlap ratios, different levels of parallax, and diverse scenarios such as indoor, outdoor, nighttime, low-light, snowy conditions, and zoom variations.
V-A2 Synthetic Dataset
@@
As shown in Fig. 7a, this test dataset was rendered using Blender. It comprises 20 distinct scenes  [^3] https://sketchfab.com [^4] https://blenderco.cn/ [^5] https://free3d.com/zh/3d-models/blender with 600 image pairs and their corresponding ground truth (GT) images. Most image pairs have large-baseline, large-parallax characteristics.
V-A3 Real-World Dataset
@@
As shown in Fig. 7b, we capture the dataset with a mobile phone. It comprises 10 distinct scenes with 682 image pairs and their corresponding GT images. We test this real-world dataset to validate the performance of PIS3R in real-world scenarios.
@@
Figure 7: Sample images from the experimental dataset. (a)Synthetic dataset rendered by Blender [blender]. (b)Real-world dataset captured by ourselves. Additionally, each scene image and its ground-truth image share the same camera extrinsic.
V-B Comparison of Image Stitching
Figure 8: Qualitative comparison of stitched images on synthetic dataset. Our PIS3R generates images with significantly more complete structures and fewer severe distortions compared to existing image stitching algorithms APAP, ELA, OBJ-GSP and UDIS++ in both datasets. Ref./Tar. Img. refers to Reference and Target Images.
Figure 9: Qualitative comparison of stitched images on real-world dataset. Our PIS3R generates images with significantly more complete structures and fewer severe distortions compared to existing image stitching algorithms APAP, ELA, OBJ-GSP and UDIS++ in both datasets. Ref./Tar. Img. refers to Reference and Target Images.
Figure 10: Qualitative comparison of stitched images on UDIS-D dataset. Even without training on the UDIS-D dataset, PIS3R remains visually comparable to various state-of-the-art methods.
All Image Pairs W/O Failed Registration Cases
Synthetic Dataset APAP 19.116 0.659 0.169 87.5% 20.063 0.691 0.146
ELA 19.813 0.670 0.158 91.5% 20.487 0.719 0.139
UDIS++ 19.858 0.690 0.139 86.2% 21.038 0.748 0.097
OBJ-GSP 16.961 0.538 0.205 82.3% 17.6733 0.563 0.1846
PIS3R 22.026 0.782 0.110 95.7% 22.351 0.796 0.102
Real World Dataset APAP 16.764 0.525 0.246 81.6% 18.047 0.572 0.207
ELA 16.983 0.547 0.245 98.2% 17.064 0.553 0.242
UDIS++ 17.390 0.554 0.204 90.3% 18.874 0.615 0.208
OBJ-GSP 15.3099 0.3918 0.2744 73.89% 15.9051 0.4123 0.2495
PIS3R 19.674 0.644 0.180 100.0% 19.674 0.644 0.180
TABLE II: Quantitative comparison of image stitching on synthetic dataset and real world dataset. PIS3R outperforms the other three methods across PSNR, SSIM and LPIPS when all image pairs are included in the evaluation. Moreover, even after excluding registration-failed cases, PIS3R still maintains a noticeable advantage in PSNR and SSIM.
Dataset Method PIQE $\downarrow$ NIQE $\downarrow$
UDIS-D APAP 44.58 4.383
ELA 47.87 4.152
UDIS++ 43.79 4.098
OBJ-GSP 56.96 4.885
PIS3R 44.13 4.059
TABLE III: Quantitative comparison of image stitching on the UDIS-D  [udis_d_nie2021unsupervised] dataset. PIS3R performs comparably to UDIS++ and both outperform other methods.
V-B1 Evaluation Metrics
We employ two sets of evaluation metrics for different datasets.
@@ V-B1 Evaluation Metrics
$\bullet$ On Dataset with GT. The traditional image stitching paradigm consists of warping and composition stages. However, as a 3D-reconstruction-based image stitching paradigm, PIS3R does not include these steps. Therefore, overlapping regions based metrics of warping results is not suitable for PIS3R. Notably, each reference image shares the same camera extrinsics with its GT image (Fig. 7), and scene images are obtained by center-cropping GT images. Theoretically, stitched images should share the same extrinsics with reference images. Thus, we propose a novel three-step metric computation approach (Fig.4 in Appendix):
@@
We employ metrics including: PSNR, SSIM  [ssim_wang2004image], LPIPS  [lpips_zhang2018unreasonable]. Furthermore, to avoid metric inaccuracies caused by registration failures, we calculated the Registration Success Rate (RSR):
@@
$\bullet$ On Dataset without GT. Following the evaluation protocol of recent state-of-the-art methods OBJ-GSP, we adopt the Naturalness Image Quality Evaluator(NIQE) [NIQE_mittal2012making] and Perception based Image Quality Evaluator(PIQE) [PIQE_venkatanath2015blind].
V-B2 Image Stitching in Synthetic Dataset
@@
$\bullet$ Qualitative Comparison. The qualitative comparison results are shown in Fig. 8, where the reference and target images are displayed in the left-most column. APAP, ELA, OBJ-GSP and UDIS++ exhibit noticeable artifacts and geometric inconsistencies, while Our results better preserve the geometric structure of the scene. Additional cases are provided in Fig. 1 of Appendix.
$\bullet$ Quantitative Comparison. The quantitative comparison results are presented in Tab. II. Our PIS3R outperforms the other three methods across all metrics when including all data in the calculations. We still maintain a substantial lead in PSNR, SSIM after excluding failed registration cases. What is noteworthy is that RSR substantially indicates the method’s ability to maintain geometric consistency.
V-B3 Image Stitching in Real-World Dataset
@@
$\bullet$ Qualitative Comparison. The qualitative comparison results are presented in Fig. 9. APAP, ELA, OBJ-GSP and UDIS++ exhibit noticeable artifacts and geometric inconsistencies. In contrast, our PIS3R method consistently generates images with significantly more complete structures and fewer severe distortions in these complex and challenging scenarios. Additional cases are provided in Fig. 2 of Appendix.
$\bullet$ Quantitative Comparison. As shown in Tab. II, PIS3R demonstrates superior performance across all evaluation metrics on the complete real-world dataset. Moreover, after excluding failed registration cases, it continues to achieve leading performance on all metrics.
V-B4 Image Stitching in UDIS-D Dataset
We evaluate PIS3R on the UDIS-D dataset, which primarily involves pure rotation and slight parallax.
$\bullet$ Qualitative Comparison. The qualitative comparison results are presented in Fig.  10. Even without training on the UDIS-D dataset, PIS3R remains visually comparable to various state-of-the-art methods.
$\bullet$ Quantitative Comparison. As shown in Tab. III, PIS3R performs comparably to UDIS++ and outperform APAP, ELA and OBJ-GSP, even without training on the UDIS-D dataset.
V-C Geometric Consistency Verification of Stitched Images
@@
$\bullet$ Qualitative Comparison. As shown in Fig. 6, we fed the stitched images of each method into AgiSoft MetaShape  [Agisoft2023] for reconstruction. PIS3R’s stitching results in VLP scenarios can be directly input into a Structure From Motion(sfm) framework for point cloud reconstruction, while better maintaining the geometric structure and integrity of the scene due to its excellent geometry preservation capability.
$\bullet$ Quantitative Comparison. As shown in Table IV, PIS3R achieves a significantly lower epipolar error for stitched images exhibiting VLP conditions, notably outperforming the other three approaches.
APAP ELA UDIS++ OBJ-GSP PIS3R
$d_{\text{Sam}}^{2}$(pixel) $\downarrow$ 22.78 38.97 12.94 16.43 4.12
TABLE IV: Sampson epipolar error of stitching image on synthetic dataset. PIS3R achieves a lower epipolar error for stitched images exhibiting very large parallax conditions.
V-D Ablation studies
All image pairs
Dataset Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Synthetic Dataset Dust3R 21.872 0.769 0.134
Mast3R 21.814 0.780 0.126
VGGT 22.754 0.794 0.121
Real World Dataset Dust3R 18.134 0.617 0.239
Mast3R 18.965 0.636 0.221
VGGT 20.134 0.652 0.207
TABLE V: Ablation studies of feed-back deep 3D reconstruction models on synthetic and real world dataset. We compare the quality of reprojected images without refine. VGGT achieves the best reprojection results in our task.
All image pairs
Dataset Model PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Synthetic Dataset w/o Refine 22.754 0.794 0.121
DiffIR 21.361 0.740 0.120
DDRM 20.894 0.706 0.132
RDDM 22.026 0.782 0.110
Real World Dataset w/o Refine 20.134 0.652 0.207
DiffIR 18.744 0.624 0.193
DDRM 18.697 0.616 0.210
RDDM 19.674 0.644 0.180
TABLE VI: Ablation studies of diffusion models on synthetic and real world dSataset. Refining with RDDM improves the perceptual quality of the stitched images (as measured by LPIPS), while having only a negligible impact on pixel-level fidelity metrics such as PSNR and SSIM.
Figure 11: Qualitative comparison of different feed-back deep 3D reoconstruction models including VGGT [wang2025vggt], MAST3R [Mast3r_leroy2024grounding] and DUST3R [dust3r_wang2024dust3r]. VGGT demonstrates visually superior reconstruction results.
V-D1 Comparison of Different 3D Reconstruction Models
We compare three feed-forward deep 3D reconstruction models: Dust3R [dust3r_wang2024dust3r], VGGT [wang2025vggt], and Mast3R [Mast3r_leroy2024grounding].
$\bullet$ Qualitative Comparison. As shown in Fig. 5, VGGT demonstrates visually superior reconstruction results.
$\bullet$ Quantitative Comparison. As shown in Tab. V, we compare the quality of reprojected images without refinement. VGGT achieves the best reprojection results in our task
V-D2 Comparison of Different Diffusion Models
We compared three image restoration diffusion models: DiffIR  [diffir_xia2023diffir] [^6] https://github.com/Zj-BinXia/DiffIR, DDRM [ddrm_kawar2022denoising] [^7] https://github.com/bahjat-kawar/ddrm, and RDDM  [liu2024residual] [^8] https://github.com/nachifur/RDDM. We train DiffIR and RDDM on 3DIS-D dataset(as illustrated in Sec. III-E) and present comparative results of the restoration performance. Notably, DDRM does not rely on task-specific training and can be applied directly. RDDM can effectively remove noise while preserving geometric structures.
$\bullet$ Qualitative Comparison. As shown in Fig. 5, RDDM demonstrates excellent optimization results due to its unique diffusion mechanism.
$\bullet$ Quantitative Comparison. As shown in Tab. VI, we quantify the metric differences between reprojected images processed and unprocessed by the diffusion model. Refining with RDDM improves the perceptual quality of the stitched images, while having only a negligible impact on pixel-level fidelity metrics such as PSNR and SSIM. This trade-off is thus deemed both acceptable and necessary.
V-E Running time
APAP1 ELA1 UDIS++2 OBJ-GSP1 PIS3R2
Time(s) $\downarrow$ 1.2 6.1 2.7 1.6 5.8
TABLE VII: Runtime of PIS3R on image stitching. 1: tested with Intel i9-14900KF 3.20GHZ CPU; 2: tested with NVIDIA RTX 4090 GPU.
As shown in Tab. VII, since we do not have the warp and composition steps, we compare the average time required to stitch a complete image, excluding the time spent loading the model. PIS3R takes 5.8 seconds to stitch an image, including 1.1 seconds for reconstruction, 3.5 seconds for reprojection, and 1.2 seconds for diffusion model refinement.
VI Conclusion